an invested adviser invested $14000 in two accounts. one investment earned 8% annual simple interest, and the other investment earned 6.5% annual simple interest. the amount of interest earned for 1 year was $1024. how much was invested in each account?
Answer by checkley71(8403)

.08*x+.065(14000-x)=1024
.08x+910-.065x=1024
.015x=1024-910
.015x=114
x=114/.015
x=7600 invested @ 8%
14,000-7,600=6,400 invested @ 6.5%
proof
.08*7600+.065*6400=1024
608+416=1024
1024=1024